OpenStudy (anonymous):

Determine all values of c so that the lines cy=4x=9 and cy-5x=10 are perpendicular. ( do I try to get each letter its own answer for example c=--, y=--, x=--? how would I do that if I do?)

OpenStudy (aum):

Are you sure the first equation is: cy=4x=9 ?

OpenStudy (anonymous):

sorry typo cy+4x=9

OpenStudy (aum):

Find the slope of the line cy + 4x = 9. Put it in the slope-intercept form: y = mx + b.

OpenStudy (aum):

cy + 4x = 9. subtract 4x: cy = -4x + 9 divide by c: y = (-4/c)x + 9/c Slope of this line is -4/c Find the slope of the line cy-5x=10 (use similar method as shown above). If the lines are perpendicular the product of their slopes = -1. Equate the product of the two slopes found above, equate it to -1 and solve for c.